OpenStudy (anonymous):

subject: formulas;dont know formulas. plz help. The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle. Find the width of the rectangle. 6 in. 6.5 in. 7.5 in.

OpenStudy (anonymous):

that's hard

OpenStudy (anonymous):

yeah i know math aint my subjec english is

OpenStudy (anonymous):

I can sepll good

OpenStudy (anonymous):

i can tell

OpenStudy (anonymous):

sory I ment spell

OpenStudy (anonymous):

Hello there, I can help you with this question. :)

OpenStudy (anonymous):

god

OpenStudy (anonymous):

thanks abunch

OpenStudy (anonymous):

sory iment good

OpenStudy (anonymous):

You are welcome. Well, I take it you are completely stuck on this problem, so I will help you. :) PLease allow me a few minutes to scratch this out, okay?

OpenStudy (anonymous):

yes ma'am

OpenStudy (anonymous):

I love inglish

OpenStudy (anonymous):

l*w right

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so where does The length of a rectangle is 4 inches more than its width come in

OpenStudy (anonymous):

You know what? The thing I was going to have you do would probably be a little more complicated. Okay, let's try something else. You know how rectangles have two long sides that match each other (lengths) and two short sides that match each other (widths)?

OpenStudy (anonymous):

yes so do leanth *2 and whith * 2 right

OpenStudy (anonymous):

Not necessarily... We're going to use those so we don't have to do so much math. Our perimeter is given as 34 inches. So if we divided that number by four, we'd get...... ?

OpenStudy (anonymous):

8.5

OpenStudy (anonymous):

do you set it up like l=w+4=34

OpenStudy (anonymous):

Sorry, had to feed the dog. :) Okay, perfect. This means that our answer must be less than or equal to 8.5 SO let's get to work at the formula you are stuck on. To find the width, you will need this formula: 2w + 2l = Perimeter And we have to substitute the whole "four inches longer than the width" thing into the equation. So if our width is "x" (we don't know what it is yet), we'll have 2w + 2(2x + 4) = 34 as our equation. Does this make sense? I can re-explain if you need me to.

OpenStudy (anonymous):

somewhat makes sence

OpenStudy (anonymous):

i see it know u can go on

OpenStudy (anonymous):

Okay, so the next step is to solve for x, so we can re-write the equation without the 2w. 2(2x + 4) = 34 Can you solve that?

OpenStudy (anonymous):

4x+8=34

OpenStudy (anonymous):

Perfect! Can you solve that equation, too?

OpenStudy (anonymous):

answer is 6.5

OpenStudy (anonymous):

You have hit the answer on the nose! Great work! :) As an extension, do you know what the length is, now that you have found the width?

OpenStudy (anonymous):

26

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

thank you so much how do i give medals

OpenStudy (anonymous):

I see why you got 26, because you multiplied 6.5 by 4; but you should have added 4 to 6.5, because the question says the length is four inches longer. :) Well, to give someone a medal, you click "best response' which is next to their answer.

OpenStudy (anonymous):

You are very welcome! :D

OpenStudy (anonymous):

thank so much

OpenStudy (anonymous):

I am glad to help. So you understand it better now?

OpenStudy (anonymous):

yes alot better thanks to you

OpenStudy (anonymous):

Great! Well, you have a great day! :)

OpenStudy (anonymous):

u 2